OpenStudy (anonymous):

Convert 50 mi/h to feet per minute.

OpenStudy (karatechopper):

5280 feet in mile.. 60 mins in an hour..

OpenStudy (mertsj):

Remember that units cancel just like numbers so set it up like this:

OpenStudy (mathmate):

1 mi/h = 5280 ft/60 min =5280/60 ft/min

OpenStudy (anonymous):

a)4,400 ft/min b)3,000 ft/min c)1,200 ft/min d)6,636 ft/min

OpenStudy (karatechopper):

50/5280=feet 60/60= min

OpenStudy (mertsj):